OpenStudy (anonymous):

Solve for C? ab-c=9 A.-ab- 9 B. 9 +ab C. ab- 9 HelP!

ganeshie8 (ganeshie8):

ohk..

ganeshie8 (ganeshie8):

lol ok :)

OpenStudy (anonymous):

A

OpenStudy (uri):

@Zofia1 Stop giving her the ansur!

ganeshie8 (ganeshie8):

since ive started this q, il try to work this... see if it makes sense :- ab-c=9 add c both sides ab-c = 9 +c +c ab = 9 + c

OpenStudy (anonymous):

Why not, lol

ganeshie8 (ganeshie8):

subtract 9 both sides and see wat u get

OpenStudy (anonymous):

uha.... GOT IT

ganeshie8 (ganeshie8):

A is incorrect btw :)

OpenStudy (anonymous):

B Agh this keyboard.

OpenStudy (uri):

i know an easy way to solve this and @Zofia1 wrong :P

OpenStudy (anonymous):

That's how I was taught@uri!!

OpenStudy (anonymous):

Im pretty shure its correct..

OpenStudy (anonymous):

ok now ya going to fight about my question

OpenStudy (anonymous):

Nope im out peace~

OpenStudy (uri):

@juliesandovaljimenez No? We want you to do it.

OpenStudy (anonymous):

i am

OpenStudy (uri):

So do it?

OpenStudy (anonymous):

i am @uri